In an isosceles trapezoid the 5- cm altitude drawn from the endpoint of the shorter base to

the larger base divides the larger base in segments of 10 cm and 25 cm long. Compute for

the area of the trapezoid.


1
Expert's answer
2022-03-31T14:18:13-0400

The altitude divides the longer base into 10cm and 25cm. If we take away 10cm from the 25cm on the other side, we're left with 15cm. This corresponds to the length of the shorter base( they are parallel to each other)

Hence, a = 15cm, b= (10+25)cm = 35 cm, h = 5cm

A = 0.5(a+b)h = 0.5(15+35)*5 = 0.5*40*5 = 100cm²